# OmniArch: Building the Foundation Model for Scientific Computing

### v2

@@ 1 Introduction
Developing robust neural surrogate models for temporal partial differential equations (PDEs) is crucial for various scientific and engineering applications, including aircraft design, weather forecasting, and semiconductor manufacturing physic_design ; fourcaset_net . These PDEs describe spatial-temporal dynamic systems that are foundational to these industries. Traditional scientific computing methods, such as Finite Element Methods (FEMs) and Finite Volume Methods (FVMs) FEM , require extensive handcrafted coding and are computationally intensive, even on state-of-the-art High-Performance Computing (HPC) clusters. To expedite PDE solving, pioneers have explored the construction of neural operators that learn mappings between function spaces, offering the potential to generalize across different discretizations. For the requisite precision, neural operators are often enhanced with physics-informed normalization techniques, such as customized loss functions derived from the governing physical equations PINNs .
The primary limitation of neural operator methods lies in their case-specific design, restricting their application scope and hindering broad transferability across diverse physical systems. Recent efforts aim to enhance the transferability of neural operators by developing foundational models that leverage advancements in learning strategies, architectural design, and data curation. In terms of learning, the pre-train and fine-tune paradigm, proven effective for Fourier Neural Operator (FNO) models fno_foundation , has been adapted to PDE contexts. Additionally, Lie group-based self-supervised learning (Lie-SSL) Lie_Symmetries introduces physics-constrained transformations for PDEs, primarily addressing inverse problems. Architecturally, innovations like ICON_LM ICON_LM , and PITT PITT incorporate language model principles to enhance neural operator learning, enabling generalization through equation captions. The Factformer Scalable_transformer introduces a scalable transformer for multi-dimensional PDE data, with the Multi-Physics Pre-training (MPP) MPP further extending this approach to 2D data pre-training. From a data-centric viewpoint, resources such as PDEBench PDEbench and PDEArena pdeareana offer well-structured datasets that facilitate pre-training and the establishment of rigorous benchmarks.
@@
Figure 1: OmniArch achieves state-of-the-art performance (nRMSE Loss) on a broad range of tasks. The baselines include the task-specific expert models PINNs ; unet ; FNO and the pre-trained models MPP ; pdeformer ; swin .
@@
While attempting unified learning of multiple PDE solvers in a single model, multi-scale and multi-physics challenges persist. The above surrogate models, often constrained by the fixed mapping grid (MPP, Lie-SSL, ICON_LM) and single-time step observation window (MPP, Facyformer, PITT), struggle with flexible spatial grid input and long-sequence roll-out predictions.
@@ 2 Related Works
Learned PDE Solvers. Deep Learning for solving PDEs has been a recent focal point of research deepxde ; physicsml , including physics-informed methods PINNs , GNN-based techniques  velivckovic2017graph ; pfaff2020learning , and neural operator models like DeepONet  deeponet and FNO FNO . While effective, these models often require task-specific training and struggle with generalization. ICON_LM icon , MPP MPP , PDEformer-1 pdeformer aim to generalize across diverse physical systems but limit to a single dimension.
Foundation Models. The Foundation Models bert ; gpt1 ; gpt2 ; gpt3 ; llama ; CLIP have emerged as pivotal elements in the field of natural language processing, computer vision, and cross-modal tasks. After large-scale pre-trained with the transformer backbone, they serve as the bedrock for a multitude of downstream tasks by fine-tuning fine-tuning or in-context learning in-context-learning . Recently, they have shown promise in scientific fields, exemplified by FourcastNet fourcaset_net for weather forecasting, OpenLAM openlam for chemistry, and HyenaDNA HyenaDNA for biomedical tasks. However, applying foundation models to scientific computing, particularly PDE solving, remains an emerging and pioneering area.
@@ 3 Method
The foundation models bert ; gpt1 ; gpt2 ; gpt3 ; llama ; CLIP have shown significant success with broad generation to various inputs and downstream tasks. Building a similar model for scientific computing should require addressing dynamic and complex physical systems and learning intrinsic laws from wild physical phenomena. We highlight the major challenges as three-fold:
@@ 3.1.1 Encoder/Decoder in Fourier Domain
The multi-scale challenge needs proper representation of inputs from different dimensions, varying grid resolutions, and shapes. Inspired by the Fourier transforms FFT convert the sequential signals into frequency components, we re-organize the multi-scale inputs in the spatial domain into the multi-component ones in the frequency domain. The traditional pipeline includes convolutional encoders traditional_conv_encoder , which capture the local features in separated dimensions while the global information exchange happens at the channels’ explicit mixing. The results of Fourier transforms are complex coefficients that measure the magnitude and phase of decomposed periodic components and the global information is naturally weighted, which also applies to the complex boundary conditions and heterogeneous grids. Based on that, we further introduce the filter-like components selecting mechanism that distinguishes the high-frequency (detailed variations) and low-frequency (overall trends) ones in physical inputs, which may maintain different patterns and distribution ratios among the local and global representation. Thus, we can build a universal representation with different resolutions and grid shapes in one flexible network architecture.
@@
$u^{\text{pred}}(x^{(d)},t+1)=\mathbf{\Psi}^{{}^{\prime}}(~{}\text{IFFT}(~{}\text{Zero-Padding}(~{}[\hat{u}^{\text{pred}}_{K}(1,t),\ldots,\hat{u}^{\text{pred}}_{K}(K,t)]~{}))~{}).$ (2)
@@ 3.1.2 Transformer as an Integral Neural Operator
$u^{\text{pred}}(x,t+1)=\Sigma^{t}_{i=1}\alpha_{i,t}\cdot u(x,i)=u(x,t)+\Sigma^{t-1}_{i=1}\alpha_{i,t-1}u(x,i),$ (4)
where $\alpha_{i,t}$ refer to the attention weights. Both approaches update based on previous time steps, with the attention mechanism acting as a neural surrogate surrogate for the integral operator $f$.
@@
where $i$ and $j$ represent the $i$-th and $j$-th tokens in the sequence, and $\left\lfloor\frac{i}{C}\right\rfloor$ represents the time step. The input sequence then passes through multiple shared Transformer blocks, outputting the shifted right predicted feature sequence for each time step $\{\mathbf{\hat{Z}}_{t}\}_{t=2}^{T+1}$:
@@
$L^{u}_{\text{sim}}=\frac{1}{|B|}\sqrt{\Sigma_{(x,t)\in B}\left(\frac{u^{\text{pred}}(x,t)-u(x,t)}{\sigma_{u}}\right)^{2}},\qquad L_{\text{sim}}=\frac{1}{C}\Sigma_{j\in C}L^{j}_{\text{sim}}.$ (8)
@@ 3.2 Fine-tuning OmniArch: Enabling Physics-Informed Learning via Equation Supervision
The PDE equations are natural and intuitive ‘supervision’ methods for real-world physical phenomena. To perform the physical alignment, we incorporate the PDE-Aligner to achieve physics-informed learning. Unlike the pre-training stage, the OmniArch is designed to comply with specific physical laws during fine-tuning. As illustrated in Figure 3 (left), the PDE-Aligner employs a contrastive learning paradigm in the frequency domain. It compares the dynamic system’s semantics with statistical characters of the frequency domain, where the dynamical system descriptions, namely equations, boundaries, initial conditions, and other physical priors, are encoded into a representation $E_{\text{text}}(\mathcal{P})$. Then, we acquire the initial state $u(x,t_{0})$ and the current state $u(x,t_{i})$ of the physical field, applying the Fourier Encoder to obtain their $k$-th frequency domain representations $\hat{u}_{K}(k,t_{i})$ and $\hat{u}_{K}(k,t_{0})$. We use the phase difference $\Delta\phi=(\hat{u}_{K}(k,t_{i})\cdot\hat{u}_{K}^{\ast}(k,t_{0}))/(|\hat{u}_{K}(k,t_{i})||\hat{u}_{K}(k,t_{0})|)$ and magnitude ratio $R={|\hat{u}_{K}(k,t_{i})|}/{|\hat{u}_{K}(k,t_{0})|}$. Thus, we have the alignment loss function as:
$L_{\text{Align}}=L_{\textrm{eq}}+\lambda L_{\textrm{E}},\qquad L_{\textrm{eq}}=\mathcal{S}(~{}E_{\text{text}}(\mathcal{P}),\mathbf{\Psi}[\Delta\phi,R]^{\top}),\qquad L_{\textrm{E}}=|\sum\nolimits_{K}R-1|.$ (9)
@@
where $\lambda$ is a hyperparameter balancing the energy conservation loss. By minimizing the alignment loss function $L_{\text{Align}}$, the PDE-Aligner aligns the changes in the physical field with the textual descriptions within the constraints of energy conservation. In the fine-tuning stage (Figure 3 Right), the pre-trained PDE-Aligner can be used to measure the alignment between the predicted physical field and the textual description of the given physical system. Recalling the $L_{\text{sim}}$ loss in Eq.(8), we reset the loss function for fine-tuning as $L_{\text{ft}}=L_{\text{sim}}-L_{\text{eq}}$.
@@ 4.1 Dataset and Baselines
Baselines. The baselines are divided into two categories: (1) Task-specific expert models, which include Physics-Informed Neural Networks (PINNs) PINNs , U-Net unet , and Fourier Neural Operator (FNO) FNO , all of which require training from scratch for each specific case (each equation/coefficient, etc.). (2) Unified pre-training models, which include PDEformer-1 pdeformer , Multiple Physics Pre-training (MPP) MPP and a SWIN-transformer swin used for the ORCA task. More details on the baselines are provided in Appendix D.
Training Details. The OmniArch model uses single-layer encoders and decoders for data of various dimensions, with the LLaMA model (trained from scratch) as the shared Transformer architecture. The PDE-Aligner employs the pre-trained Fourier encoder from OmniArch to encode physical fields and the pre-trained BERT-base-cased model to encode PDE captions. Two versions of the model are provided: Base and Large. Additional training details are available in Appendix E.
@@ 4.2 Results and Analysis
1D 2D 3D Methods CFD  Adv.  Bur.  Diff.  Reac. CFD   Reac.   SWE   Incom CFDMaxw. Baselines -Task specific Expert Models PINNs    /   0.8130  0.9450  0.2200  0.2140    /    1.6000  0.0170    /    /        / U-Net 2.6700  0.7760  0.3201  0.1507  0.0026 1.0700  0.8401  0.0830  1.1200 0.7989  0.2999 FNO 1.4100  0.0091  0.0174  0.0017  0.0005 0.2060  0.1203  0.0044  0.2574 0.3052  0.1906 Baselines - Unified Pre-training and Fine-tuning PDEformer-1    –   0.0043   0.0095  –      0.0009 –     –    –     – –     – ORCA-SWIN-B –     –    –     –     –    /      0.8201  0.0062    / –     – MPP-AVIT-B –     –    –     –     – 0.0227  0.0106  0.0024    / –     – MPP-AVIT-L –     –    –     –     – 0.0178  0.0098  0.0022    / –     – Full Pre-Training on 1D,2D,3D Data OmniArch-B(Ours) 0.0340  0.0238  0.0089  0.0020  0.0006 0.0196  0.0158  0.0016  0.1726 0.5209  0.2834 OmniArch-L(Ours) 0.0250  0.0182  0.0063  0.0015  0.0004 0.0148  0.0105  0.0014  0.1494 0.4531   0.2268  + PDE-Aligner Fine-tuning OmniArch-B(Ours) 0.0302  0.0201  0.0071  0.0017  0.0003 0.0153  0.0102  0.0015  0.0955 0.4032  0.1813 OmniArch-L(Ours) 0.0200  0.0041  0.0032  0.0006  0.0002 0.0125  0.0084  0.0012  0.0827 0.3723  0.1671  std. $\pm$ 0.0031  0.0012  0.0004  0.0001  0.0001 0.0017  0.0004  0.0003  0.0023 0.0443  0.0197 Improvement $\uparrow$ 98.70%   4.65%   66.32%   64.75%   60.00% 29.70%  14.28%  45.45%  67.87%  –  12.32%
@@
Multi-Physics Results. (1) Compared with Task-specific Expert Models. PINNs, U-Net, and FNO require training from scratch for each specific equation or coefficient. While FNO shows strong performance, PINNs and U-Net struggle with convergence and accuracy in some cases (Like the CFD-1D, and CFD-2D). (2) Compared with Unified Pre-training Models. PDEformer-1 exhibits proficiency in specific 1D equations but fails to generalize beyond its formulation structure. MPP and ORCA-SWIN leverage 2D pre-training and fine-tuning, improving generalization, yet their effectiveness remains constrained by the diversity of the pre-training data. (3) OmniArch Performance. OmniArch, pre-trained on 1D, 2D, and 3D data, demonstrates superior performance across all evaluated datasets. Both the base (B) and large (L) versions of OmniArch outperform existing models, validating its robustness in multi-physics contexts. (4) PDE-Aligner Fine-tuning. Fine-tuning with PDE-Aligner significantly enhances OmniArch’s accuracy, particularly for complex datasets. This step utilizes a pre-trained Fourier encoder and BERT-base-cased model, ensuring precise alignment between physical fields and PDE descriptions. OmniArch demonstrates substantial performance gains over baselines, with up to 98.70% improvement on CFD-1D and notable enhancements across other PDEs. These results corroborate its superior multi-physics generalization driven by comprehensive pre-training and effective physics-language alignment through PDE-Aligner.
@@
Multi-scale Results. In Figure 4, we present the multi-scale inference performance of OmniArch-Base and OmniArch-Large on the 2D Incom. Dataset. Due to the frequency truncation capability of the Fourier Encoder, OmniArch can handle inputs of varying grid sizes without requiring re-training. In the red-shaded area, the nRMSE decreases as the grid size becomes smaller. Conversely, in the blue-shaded area, the nRMSE slightly increases. However, even with a grid size of 512, the maximum nRMSE remains below 0.2. In the rollout settings, a grid size of 256 sometimes leads to better or comparable performance to a grid size of 128. Additional visualizations are provided in Appendix H.5.
Table 2: The Performance on Zero-shot PDEs.
@@
Zero-shot Performance. Our examination of 2D PDE predictions, as illustrated in Figure 5, reveals that OmniArch effectively captures both low- and high-frequency patterns even in zero-shot scenarios, surpassing former 2D models like MPP. MPP often misses key features, leading to erroneous representations of the primary physics and failed rollouts. Details on the zero-shot dataset, including shock wave, KH, and OTVortex phenomena, are provided in Appendix C.2.
@@
As shown in Table 4, nRMSE scores indicate that all models, except OmniArch, tend to underperform in zero-shot transfer. This suggests that OmniArch’s use of Fourier Encoders and unified training approach enhances its ability to generalize across different PDEs. By leveraging flexible grid inputs and dynamic observation windows during pre-training, OmniArch effectively captures the underlying physics of the observed field states, which may not be adequately addressed by methods adhering strictly to explicit grid and temporal dependencies.
In-Context Learning. After auto-regressively pre-trained on various dynamic systems, we observe that OmniArch could learn neural operators within the observations of several time steps, which is similar to the in-context learning in Large Language Models. Here, we define the given time-series of observations as PDE Prompt. Our approach varies the prompt length from 2 tokens (derived from a 50 time step interval) to 100 tokens (from a 1 time step interval). As indicated in Figure 7, longer prompts yield higher precision with less variance, while shorter prompts can expedite inference by up to 10 times compared to full-length prompts. More details are in Appendix H.2.

### v3

# OmniArch: Building Foundation Model For Scientific Computing

@@ 1 Introduction
Developing robust neural surrogate models for temporal partial differential equations (PDEs) is crucial for various scientific and engineering applications, including aircraft design, weather forecasting, and semiconductor manufacturing (Allen et al., 2022; Pathak et al., 2022). These PDEs describe spatial-temporal dynamic systems that are foundational to these industries. Traditional scientific computing methods, such as Finite Element Methods (FEMs) and Finite Volume Methods (FVMs) (Oden, 1989), require extensive handcrafted coding and are computationally intensive, even on state-of-the-art High-Performance Computing (HPC) clusters. To expedite PDE solving, pioneers have explored the construction of neural operators that learn mappings between function spaces, offering the potential to generalize across different discretizations. For the requisite precision, neural operators are often enhanced with physics-informed normalization techniques, such as customized loss functions derived from the governing physical equations (Raissi et al., 2019).
@@
Figure 1: OmniArch achieves state-of-the-art performance (nRMSE Loss) on 1D-2D-3D PDE tasks with single foundation model. The baselines include the task-specific expert models and the pre-trained models.
The primary limitation of neural operator methods lies in their case-specific design, restricting their application scope and hindering broad transferability across diverse physical systems. Recent efforts aim to enhance the transferability of neural operators by developing foundational models that leverage advancements in learning strategies, architectural design, and data curation (Alkin et al., 2024; Sun et al., 2024; Shen et al., ). In terms of learning, the pre-train and fine-tune paradigm, proven effective for Fourier Neural Operator (FNO) models (Subramanian et al., 2023), has been adapted to PDE contexts. Additionally, Lie group-based self-supervised learning (Lie-SSL) (Mialon et al., 2023) introduces physics-constrained transformations for PDEs, primarily addressing inverse problems. Architecturally, innovations like ICON_LM (Yang et al., 2023b), and PITT (Lorsung et al., 2023) incorporate language model principles to enhance neural operator learning, enabling generalization through equation captions. The Factformer (Li et al., 2023) introduces a scalable transformer for multi-dimensional PDE data, with the Multi-Physics Pre-training (MPP) (McCabe et al., 2023), Poseidon (Herde et al., 2024) and DPOT (Hao et al., 2024) further extending this approach to 2D data pre-training. From a data-centric viewpoint, resources such as PDEBench (Takamoto et al., 2022), PDEArena (Gupta & Brandstetter, 2022) and The-Well (Ohana et al., 2024) offer well-structured datasets that facilitate pre-training and the establishment of rigorous benchmarks.
@@
While attempting unified learning of multiple PDE solvers in a single model, multi-scale and multi-physics challenges persist. The above surrogate models, often constrained by the fixed mapping grid (MPP, Lie-SSL, ICON_LM) and single-time step observation window (MPP, Factformer, PITT, Poseidon), struggle with flexible spatial grid input and long-sequence roll-out predictions.
@@ 2 Related Works
Learned PDE Solvers. Deep Learning for solving PDEs has been a recent focal point of research (Lu et al., 2021b; Karniadakis et al., 2021), including physics-informed methods (Raissi et al., 2019), GNN-based techniques  (Veličković et al., 2017; Pfaff et al., 2020), and neural operator models like DeepONet  (Lu et al., 2021a) and FNO (Li et al., 2020). While effective, these models often require task-specific training and struggle with generalization. ICON_LM (Yang et al., 2023a), MPP (McCabe et al., 2023), PDEformer-1 (Ye et al., 2024) aim to generalize across diverse physical systems but limit to a single dimension.
Foundation Models for Science. The Foundation Models (Devlin et al., 2018; Brown et al., 2020; Radford et al., 2019, 2018; Touvron et al., 2023; Radford et al., 2021) have emerged as pivotal elements in the field of natural language processing, computer vision, and cross-modal tasks. After large-scale pre-trained with the transformer backbone, they serve as the bedrock for a multitude of downstream tasks by fine-tuning (Zhang et al., 2023) or in-context learning (Li, 2023). Recently, they have shown promise in scientific fields, exemplified by FourcastNet (Pathak et al., 2022) for weather forecasting, OpenLAM (Zhang et al., 2022) for chemistry, and HyenaDNA (Nguyen et al., 2023) for biomedical tasks. However, applying foundation models to scientific computing, particularly PDE solving, remains an emerging and pioneering area.
@@ 3 Method
The foundation models (Devlin et al., 2018; Brown et al., 2020; Radford et al., 2019, 2018; Touvron et al., 2023; Radford et al., 2021) have shown significant success with broad generation to various inputs and downstream tasks. Building a similar model for scientific computing should require addressing dynamic and complex physical systems and learning intrinsic laws from wild physical phenomena. We highlight the major challenges as three-fold:
@@ 3.1.1 Encoder/Decoder in Fourier Domain
The multi-scale challenge needs proper representation of inputs from different dimensions, varying grid resolutions, and shapes. Inspired by the Fourier transforms (Brigham, 1988) convert the sequential signals into frequency components, we re-organize the multi-scale inputs in the spatial domain into the multi-component ones in the frequency domain. The traditional pipeline includes convolutional encoders (Raonic et al., 2023), which capture the local features in separated dimensions while the global information exchange happens at the channels’ explicit mixing. The results of Fourier transforms are complex coefficients that measure the magnitude and phase of decomposed periodic components and the global information is naturally weighted, which also applies to the complex boundary conditions and heterogeneous grids. Based on that, we further introduce the filter-like components selecting mechanism that distinguishes the high-frequency (detailed variations) and low-frequency (overall trends) ones in physical inputs, which may maintain different patterns and distribution ratios among the local and global representation. Thus, we can build a universal representation with different resolutions and grid shapes in one flexible network architecture.
@@
$\begin{split}&u^{\text{pred}}(x^{(d)},t+1)=\\ &\mathbf{\Psi}^{{}^{\prime}}(~{}\text{IFFT}(~{}\text{Zero-Padding}(~{}[\hat{u}^{\text{pred}}_{K}(1,t),\ldots,\hat{u}^{\text{pred}}_{K}(K,t)]~{}))~{}).\end{split}$ (2)
@@ 3.1.2 Transformer as an Integral Neural Operator
$u^{\text{pred}}(x,t+1)$ $=\Sigma^{t}_{i=1}\alpha_{i,t}\cdot u(x,i)$ (4)
$=u(x,t)+\Sigma^{t-1}_{i=1}\alpha_{i,t-1}u(x,i),$
where $\alpha_{i,t}$ refer to the attention weights. Both approaches update based on previous time steps, with the attention mechanism acting as a neural surrogate (Sun et al., 2020) for the integral operator $f$.
@@
where $i$ and $j$ represent the $i$-th and $j$-th tokens in the sequence, and $\left\lfloor\frac{i}{C}\right\rfloor$ represents the time step. Unlike standard causal masking that enforces strict sequential dependencies, our Temporal Mask enables all physical quantities within the same timestep to attend to each other, addressing the fundamental coupling inherent in multi-physics systems. Specifically, for a system with $C$ physical quantities at each timestep, tokens $\{i,i+1,...,i+C-1\}$ corresponding to timestep $t$ have full visibility of each other (intra-timestep attention), while maintaining causal relationships across timesteps (inter-timestep attention). This hierarchical attention pattern ensures that coupled physical quantities—such as velocity and pressure in fluid dynamics—can jointly evolve while respecting temporal causality. The design is particularly crucial for systems where physical variables must satisfy simultaneous constraints (e.g., continuity equations in Navier-Stokes) that cannot be properly modeled through sequential token processing.
The input sequence then passes through multiple shared Transformer blocks, outputting the shifted right predicted feature sequence for each time step $\{\mathbf{\hat{Z}}_{t}\}_{t=2}^{T+1}$:
@@
$L^{u}_{\text{sim}}$ $=\frac{1}{|B|}\sqrt{\Sigma_{(x,t)\in B}\left(\frac{u^{\text{pred}}(x,t)-u(x,t)}{\sigma_{u}}\right)^{2}},$ (8)
$L_{\text{sim}}$ $=\frac{1}{C}\Sigma_{j\in C}L^{j}_{\text{sim}}.$
@@ 3.2 Fine-tuning OmniArch: Enabling Physics-Informed Learning via Equation Supervision
1D 2D 3D Methods CFD  Adv.  Bur.  Diff.  Reac. CFD   Reac.   SWE   Incom CFDMaxw. Baselines -Task specific Expert Models PINNs    /   0.8130  0.9450  0.2200  0.2140    /    1.6000  0.0170    /    /        / U-Net 2.6700  0.7760  0.3201  0.1507  0.0026 1.0700  0.8401  0.0830  1.1200 0.7989  0.2999 FNO 1.4100  0.0091  0.0174  0.0017  0.0005 0.2060  0.1203  0.0044  0.2574 0.3052  0.1906 Baselines - Unified Pre-training and Fine-tuning PDEformer-1    –   0.0043   0.0095  –      0.0009 –     –    –     – –     – ORCA-SWIN-B –     –    –     –     –    /      0.8201  0.0062    / –     – MPP-AVIT-B –     –    –     –     – 0.0227  0.0106  0.0024    / –     – MPP-AVIT-L –     –    –     –     – 0.0178  0.0098  0.0022    / –     – Poseidon-L –     –    –     –     – 0.1079  0.0949  0.0243    – –     – DPOT-L –     –    –     –     – 0.0112  0.0263  0.0451    – 0.4321     – Full Pre-Training on 1D,2D,3D Data OmniArch-B(Ours) 0.0340  0.0238  0.0089  0.0020  0.0006 0.0196  0.0158  0.0016  0.1726 0.5209  0.2834 OmniArch-L(Ours) 0.0250  0.0182  0.0063  0.0015  0.0004 0.0148  0.0105  0.0014  0.1494 0.4531   0.2268  + PDE-Aligner Fine-tuning OmniArch-B(Ours) 0.0302  0.0201  0.0071  0.0017  0.0003 0.0153  0.0102  0.0015  0.0955 0.4032  0.1813 OmniArch-L(Ours) 0.0200  0.0041  0.0032  0.0006  0.0002 0.0125  0.0084  0.0012  0.0827 0.3723  0.1671  std. $\pm$ 0.0031  0.0012  0.0004  0.0001  0.0001 0.0017  0.0004  0.0003  0.0023 0.0443  0.0197 Improvement $\uparrow$ 98.70%   4.65%   66.32%   64.75%   60.00%  –  14.28%  45.45%  67.87%  –  12.32%
@@
The PDE equations are natural and intuitive ‘supervision’ methods for real-world physical phenomena. To perform the physical alignment, we incorporate the PDE-Aligner to achieve physics-informed learning. Unlike the pre-training stage, the OmniArch is designed to comply with specific physical laws during fine-tuning. As illustrated in Figure 3 (left), the PDE-Aligner employs a contrastive learning paradigm in the frequency domain.
The key insight is that physical evolution manifests distinctively in frequency space—conservation laws constrain energy distribution across modes, while different PDEs exhibit characteristic spectral signatures. By operating in this domain, PDE-Aligner captures these fundamental patterns more effectively than spatial approaches. It compares the dynamic system’s semantics with statistical characters of the frequency domain, where the dynamical system descriptions, namely equations, boundaries, initial conditions, and other physical priors, are encoded into a representation $E_{\text{text}}(\mathcal{P})$.
To characterize physical evolution, we acquire the initial state $u(x,t_{0})$ and the current state $u(x,t_{i})$ of the physical field, applying the Fourier Encoder to obtain their $k$-th frequency domain representations $\hat{u}_{K}(k,t_{i})$ and $\hat{u}_{K}(k,t_{0})$. The phase difference $\Delta\phi=(\hat{u}_{K}(k,t_{i})\cdot\hat{u}_{K}^{\ast}(k,t_{0}))/(|\hat{u}_{K}(k,t_{i})||\hat{u}_{K}(k,t_{0})|)$ captures wave propagation and dispersion characteristics, while the magnitude ratio $R={|\hat{u}_{K}(k,t_{i})|}/{|\hat{u}_{K}(k,t_{0})|}$ quantifies energy transfer across scales—both serving as physics-aware fingerprints of the underlying PDE. Thus, we have the alignment loss function as:
$L_{\text{Align}}=L_{\textrm{eq}}+\lambda L_{\textrm{E}},$ (9)
$L_{\textrm{eq}}=\mathcal{S}(~{}E_{\text{text}}(\mathcal{P}),\mathbf{\Psi}[\Delta\phi,R]^{\top}),$
$L_{\textrm{E}}=|\sum\nolimits_{K}R-1|.$
@@
where $\lambda$ is a hyperparameter balancing the energy conservation loss. The energy term $L_{\textrm{E}}$ enforces Parseval’s theorem, ensuring physical consistency in the frequency domain. By minimizing the alignment loss function $L_{\text{Align}}$, the PDE-Aligner aligns the changes in the physical field with the textual descriptions within the constraints of energy conservation.
In the fine-tuning stage (Figure 3 Right), the pre-trained PDE-Aligner serves as a physics-aware discriminator, helping OmniArch distinguish between different physical regimes encountered during pre-training. The fine-tuning loss $L_{\text{ft}}=L_{\text{sim}}-L_{\text{eq}}$ encourages predictions that are both accurate (via $L_{\text{sim}}$) and physically consistent with the specified PDE system (via $L_{\text{eq}}$), effectively steering the model toward the correct physical behavior among many learned dynamics.
@@ 4.1 Dataset and Baselines
Baselines. The baselines are divided into two categories: (1) Task-specific expert models, which include Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019), U-Net (Ronneberger et al., 2015), and Fourier Neural Operator (FNO) (Li et al., 2020), all of which require training from scratch for each specific case (each equation/coefficient, etc.). (2) Unified pre-training models, which include PDEformer-1 (Ye et al., 2024), Multiple Physics Pre-training (MPP) (McCabe et al., 2023), SWIN-transformer (Liu et al., 2021) used for the ORCA task, the large size pretrained checkpoint of Poseidon (Herde et al., 2024) and DPOT (Hao et al., 2024). More details on the baselines are provided in Appendix D.
Training Details. The OmniArch model uses single-layer encoders and decoders for data of various dimensions, with the LLaMA model (trained from scratch) as the shared Transformer architecture. The PDE-Aligner employs the pre-trained Fourier encoder from OmniArch to encode physical fields and the pre-trained BERT model to encode PDE captions. Additional training details are in Appendix E.
@@ 4.2 Results and Analysis
Multi-Physics Results. (1) Compared with Task-specific Expert Models. PINNs, U-Net, and FNO require training from scratch for each specific equation or coefficient. While FNO shows strong performance, PINNs and U-Net struggle with convergence and accuracy in some cases (Like the CFD-1D, and CFD-2D). (2) Compared with Unified Pre-training Models. PDEformer-1 exhibits proficiency in specific 1D equations but fails to generalize beyond its formulation structure. MPP and ORCA-SWIN leverage 2D pre-training and fine-tuning, improving generalization, yet their effectiveness remains constrained by the diversity of the pre-training data. Poseidon enables single-step inference at arbitrary timesteps, though its accuracy still leaves room for improvement. DPOT successfully transfers knowledge from 2D to 3D CFD through weight sharing, but it lacks support for 1D CFD and its performance on non-CFD physics systems requires further enhancement. (3) OmniArch Performance. OmniArch, pre-trained on 1D, 2D, and 3D data, demonstrates superior performance across all evaluated datasets. Both the base (B) and large (L) versions of OmniArch outperform existing models, validating its robustness in multi-physics contexts. To validate our architectural design choices, we conduct ablation studies on the Temporal Mask mechanism (Table 2). The results confirm that our Temporal Mask, which enables full attention among physical quantities within each timestep, significantly outperforms standard causal masking across various multi-physics systems. (4) PDE-Aligner Fine-tuning. Fine-tuning with PDE-Aligner significantly enhances OmniArch’s accuracy, particularly for complex datasets. This step utilizes a pre-trained Fourier encoder and BERT-base-cased model, ensuring precise alignment between physical fields and PDE descriptions. Table 3 quantifies the impact of PDE-Aligner across different dimensions, showing consistent improvements of over 20% compared to pre-training alone. OmniArch demonstrates substantial performance gains over baselines, with up to 98.70% improvement on CFD-1D and notable enhancements across other PDEs.
Table 2: Ablation study on masking strategies
Dataset Causal Mask No Mask Temporal Mask
2D Incom. 0.0277 0.0285 0.0227
2D CFD 0.0198 0.0205 0.0148
3D CFD 0.1842 0.1923 0.1494
Ablation Study on Masking Strategies. As illustrated in Table 2, the superiority of Temporal Mask (18-20% improvement) reveals a fundamental insight: multi-physics systems require simultaneous rather than sequential processing of coupled variables. This advantage is most pronounced in 3D CFD, where the complex interplay between five physical quantities (velocities, density, pressure) demands holistic attention patterns.
Table 3: Impact of PDE-Aligner on model performance (OmniArch-L)
Configuration 1D PDEs 2D PDEs 3D PDEs
Pre-training only 0.0103 0.0440 0.3399
Fine-tuning w/o Aligner 0.0073 0.0345 0.3432
Fine-tuning w/ Aligner 0.0056 0.0262 0.2697
Improvement 23.3% 24.1% 21.4%
Impact of PDE-Aligner. We report the impact of PDE-Aligner in Table 3, where the consistent 22% improvement across dimensions suggests that PDE-Aligner serves as more than a physics constraint—it helps OmniArch disambiguate between different physical regimes learned during pre-training. Notably, the similar improvement ratios across 1D-3D indicate that physical alignment is dimension-agnostic, validating our unified architecture design.
@@
Multi-scale Results. In Figure 4, we present the multi-scale inference performance of OmniArch-Base and OmniArch-Large on the 2D Incom. Dataset. Due to the frequency truncation capability of the Fourier Encoder, OmniArch can handle inputs of varying grid sizes without requiring re-training. In the red-shaded area, the nRMSE decreases as the grid size becomes smaller. Conversely, in the blue-shaded area, the nRMSE slightly increases. However, even with a grid size of 512, the maximum nRMSE remains below 0.2. In the rollout settings, a grid size of 256 sometimes leads to better or comparable performance to a grid size of 128. The non-monotonic relationship between grid resolution and error (red vs. blue regions in Figure 4) reveals an intriguing property of frequency-domain learning: OmniArch naturally identifies the intrinsic resolution of physical phenomena. The optimal performance at intermediate resolutions (128-256) suggests the model has learned to distinguish between meaningful physical scales and numerical artifacts. Additional visualizations are provided in Appendix H.5.
@@
Zero-shot Performance. Our examination of 2D PDE predictions, as illustrated in Figure 5, reveals that OmniArch effectively captures both low- and high-frequency patterns even in zero-shot scenarios, surpassing former 2D models like MPP. MPP often misses key features, leading to erroneous representations of the primary physics and failed rollouts. Details of zero-shot dataset are in Appendix C.2.
@@
As shown in Table 4, nRMSE scores indicate that all models, except OmniArch, tend to underperform in zero-shot transfer. This suggests that OmniArch’s use of Fourier Encoders and unified training approach enhances its ability to generalize across different PDEs. By leveraging flexible grid inputs and dynamic observation windows during pre-training, OmniArch effectively captures the underlying physics of the observed field states, which may not be adequately addressed by methods adhering strictly to explicit grid and temporal dependencies. The 4-7× error reduction compared to MPP in zero-shot scenarios (Table 4) indicates that OmniArch has learned transferable physical operators rather than dataset-specific patterns. The success on shock-dominated flows (Shock, KH)—notoriously difficult for neural methods—demonstrates that frequency-domain representations capture discontinuities more effectively than spatial approaches.
Figure 7: The confusion matrix of the PDE-Aligner classification results.
In-Context Learning. After autoregressively pre-trained on various dynamic systems, we observe that OmniArch could learn neural operators within the observations of several time steps, which is similar to the in-context learning in Large Language Models. Here, we define the given time-series of observations as PDE Prompt. Our approach varies the prompt length from 2 tokens (derived from a 50 time step interval) to 100 tokens (from a 1 time step interval). More details are in Appendix H.2.